# Soft Condorcet Optimization for Ranking of General Agents

### v2

@@ Abstract
A common way to drive progress of AI models and agents is to compare their performance on standardized benchmarks. Comparing the performance of general agents requires aggregating their individual performances across a potentially wide variety of different tasks. In this paper, we describe a novel ranking scheme inspired by social choice frameworks, called Soft Condorcet Optimization (SCO), to compute the optimal ranking of agents: the one that makes the fewest mistakes in predicting the agent comparisons in the evaluation data. This optimal ranking is the maximum likelihood estimate when evaluation data (which we view as votes) are interpreted as noisy samples from a ground truth ranking, a solution to Condorcet’s original voting system criteria. SCO ratings are maximal for Condorcet winners when they exist, which we show is not necessarily true for the classical rating system Elo. We propose three optimization algorithms to compute SCO ratings and evaluate their empirical performance. When serving as an approximation to the Kemeny-Young voting method, SCO rankings are on average 0 to 0.043 away from the optimal ranking in normalized Kendall-tau distance across 865 preference profiles from the PrefLib open ranking archive. In a simulated noisy tournament setting, SCO achieves accurate approximations to the ground truth ranking and the best among several baselines when 59% or more of the preference data is missing. Finally, SCO ranking provides the best approximation to the optimal ranking, measured on held-out test sets, in a problem containing 52,958 human players across 31,049 games of the classic seven-player game of Diplomacy.
1 Introduction
77footnotetext: Now at Meta, work done while at Google DeepMind.
@@ 1 Introduction
Progress in the field of artificial intelligence has been driven by measuring the performance of agents on common benchmarks and challenge problems. AI pioneers designed checkers-playing agents (Samuel, 1959), then strove for super-human performance chess (Campbell et al., 2002), and then in other games such as Backgammon (Tesauro, 1995), Arimaa (Syed, ), Go (Silver et al., 2016), Starcraft (Vinyals et al., 2019) and Diplomacy († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, Jacob, Komeili, Konath, Kwon, Lerer, Lewis, Miller, Mitts, Renduchintala, Roller, Rowe, Shi, Spisak, Wei, Wu, Zhang, and Zijlstra, FAIR). In machine learning, common benchmarks like the UCI data set repository allowed direct comparisons of supervised learning algorithms (Kelly et al., ). Competitions, such as ImageNet, led to breakthroughs in deep learning (Krizhevsky et al., 2012).
@@
All of these examples require comparing agents (or models). Original success stories such as DeepBlue, TD-Gammon, and AlphaGo focused on a single domain. In the past ten years, agents have become increasingly more generally capable. AlphaZero extended application of AlphaGo to chess and Shogi (Silver et al., 2018). The Arcade Learning Environment (Bellemare et al., 2013), which steered much of the agent development in deep reinforcement learning, evaluated agents across 57 different Atari games. Recently, language models have been evaluated across suites of tasks such as in HELM (Liang et al., 2022), BIG-bench (bench authors, 2023) AgentBench (Liu et al., 2023), and via a public leaderboard such as Chatbot Arena driven by human voting (Chiang et al., 2024). Answering simple questions for these generally capable agents, such as “Which is the best agent?” or “Is agent $X$ better than agent $Y$?” or “What is the relative ranking of agents $X$, $Y$, and $Z$?” become increasingly more difficult when aggregating over many different contexts: how agents are scored may vary wildly across tasks, data collected for evaluation may not be balanced evenly across tasks (or agents), and classical rating systems were simply not designed for this use case.
@@
To address these problems, recent ranking methods such as Vote’N’Rank (Rofin et al., 2023) and Voting-as-Evaluation (VasE) (Lanctot et al., 2023) use voting methods to aggregate results across tasks. Using computational social choice as a basis for ranking agents has several benefits: these systems have been well-studied and understood for a long time, they inherit consistency properties of the voting methods they are based on, they do no require score normalization, and are less sensitive to score values and agent population than game-theoretic evaluation schemes such as Nash averaging (Balduzzi et al., 2018). However, classic voting schemes, and related tournament solutions, typically assume that the data (e.g. agent comparisons) is complete. This assumption is not necessarily valid in the agent evaluation setting. For example, on the webDiplomacy web site (webDiplomacy Development Team, ) there were 52,958 unique (human) agents across only 31,049 seven-player games played between 2008 and 2019: most players have not played against most other players; hence, there are no direct comparisons between the vast majority ($\geq 99\%$) of agent pairs. While there has been research on identifying “necessary and possible winners” when there is incomplete voting or comparison data, the results are mixed (Pini et al., 2011; Xia and Conitzer, 2011; Aziz et al., 2015) and most of the findings are focused on identifying top-ranked agents, not ranking all agents as is our focus.
@@
In this paper, we introduce a new ranking scheme for general agents inspired by the interpretation of voting rules as maximum likelihood estimators (Conitzer and Sandholm, 2005). Starting with Condorcet’s original model of voting (Brandt et al., 2016, Chapter 8), Young showed that the maximum likelihood estimate (MLE) of the true ranking is the one that minimizes the sum of Kendall-tau distances to all the votes. This also corresponds to the ranking found by Kemeny’s voting rule (Kemeny, 1959). Finding this solution directly is computationally expensive and does not scale to many agents, a regime that is common in the agent evaluation setting. Like classical rating systems for agent evaluation (such as Elo (Elo, 1978) and TrueSkill (Herbrich et al., 2006)), Soft Condorcet Optimization (SCO) assigns a numerical rating (score) $\theta_{a}$ to each agent (alternative) $a$. SCO then treats these ratings as a parameter vector, the votes as a data set, and defines a differentiable loss function as the objective. The final ranking of agents is obtained by sorting the ratings.
In summary, this paper makes the following contributions:
  • The SCO ranking scheme with the following properties:
@@
    • Three optimization methods to find ratings and corresponding rankings: gradient descent applied to (i) a soft Kendall-tau distance (“sigmoid loss”), or (ii) a Fenchel-Young loss (perturbed optimization) (Berthet et al., 2020); or (iii) solving a sigmoidal program with a branch-and-bound method (Udell and Boyd, 2014).
    • Online forms that can update ratings, and thus rankings, from individual outcomes as evaluation data arrives.
    • Theorem 1, guaranteeing that the top-ranked agent by SCO ratings according to the sigmoid loss is the Condorcet winner when one exists.
  • Empirical evaluations that demonstrate the following:
    • SCO ranking using sigmoid loss solves a failure mode of classical Elo rating system which may top-rank an agent that is not a Condorcet winner even when one exists.
    • SCO can serve as an approximation to the Kemeny-Young voting method, indeed empirically finding low approximation error to the optimal ranking: on average 0 to 0.043 away in normalized Kendall-tau distance across 865 preference profiles from the PrefLib (Mattei and Walsh, 2013) repository.
    • In a noisy tournament setting with sparse data, SCO approximates the true ranking best when a large proportion (59% or more) of the data is missing.
    • SCO ratings are closer to optimal rankings than Elo and voting-as-evaluation methods on held out test sets over 31,049 human Diplomacy games played by 52,958 players.
2 Background
@@
2.1 Evaluation of General Agents
We first paraphrase several key descriptions from (Balduzzi et al., 2018; Lanctot et al., 2023). The problem of evaluating agents is that of ranking agents according to their skill. Skill can be determined in several ways. In the Agent-versus-Task (AvT) setting, agents compete individually in different tasks and compare outcomes (scores) to each other in each task (e.g., language models and the various metrics for assessing their abilities). In the Agent-versus-Agent (AvA) setting, agents directly compete against each other (e.g., online games such as chess or Diplomacy).
2.1.1 Classical Evaluation
@@
Elo has a number of positive qualities. First, it is a simple rule. Second, it can be used to estimate win rates between any two agents. Third, it can be easily employed online, i.e., to modify players’ ratings from the result of a single game. It is also a special case of logistic regression (for details, see Appendix C.2). Hence, Elo has been widely applied and is a common default choice for evaluation of agents. However, Elo has a number of well-known drawbacks (Balduzzi et al., 2018; Lanctot et al., 2023; Bertrand et al., 2023). Of particular interest is the incompatibility of Elo with the concept of a Condorcet winner from social choice theory; examples are summarized in Section 4.1.
2.1.2 Voting as Evaluation of General Agents
Another way to evaluate agents is to use social choice theory, called Voting-as-Evaluation (VasE) recently proposed in (Lanctot et al., 2023). In VasE, the alternatives correspond to general agents and votes to assessments of their performance. In the AvT setting, agents are ordered based on their performance on different tasks (such as each game in the Atari Learning Environment (Bellemare et al., 2013) or on different benchmarks for large language models (Liu et al., 2023)). In the AvA setting, agents compete directly in a multiagent environment, such as multiplayer games (like chess or poker), and each game outcome corresponds to a ranking over a subset of agents. Chatbot Arena (Zheng et al., 2023), where language models compete head-to-head to provide the best answer to the same question, is another instance of the AvA setting. Casting agent evaluation as an application of computational social choice provides the benefit of robustness in the form of Condorcet consistency, clone/composition consistency, agenda consistency, and/or population consistency depending on the choice of voting rule used to evaluate agents. However, it is unclear how well these methods perform when the data is missing or unevenly distributed, which often occurs in the agent evaluation setting. SCO is motivated similarly to VasE and, as such, will be empirically assessed mainly for agent evaluation. To make the connection to ideas from the social choice literature, we will often use voting language when discussing SCO and its use in evaluation. Relevant terminology is introduced later in the paper.
2.2 Permutation and Ranking Distances
@@
2.3 Social Choice Theory
A voting scheme is defined as $\langle A,V,f\rangle$ where $A=\{a_{1},\ldots,a_{m}\}$ is the set of alternatives (agents), $V=\{v_{1},\ldots,v_{n}\}$ is the set of voters, and $f$ is the voting rule that determines how votes are aggregated. Voters have preferences over alternatives: $a_{1}\succ_{v_{i}}a_{2}$ indicates voter $v_{i}$ strictly prefers alternative $a_{1}$ over alternative $a_{2}$, and $a_{1}\succeq_{v_{i}}a_{2}$ indicates the voter has a weak preference. These preferences induce (nonstrict) total orders over alternatives, which we denote by $\mathcal{L}$. A preference profile, $[\succeq]\in\mathcal{L}^{n}$, is a vector specifying the preferences of each voter in $V$. It can be useful to summarize the preference profile in a voter preference matrix $\mathbf{N}$ or vote margin matrix $\mathbf{M}$. The preference count $N(x,y)$, $x,y\in A$ is the number of voters in []delimited-[]succeeds-or-equals[$\succeq$][ ⪰ ] that strictly prefer $x$ to $y$. The vote margin is the difference in preference count: $\delta(x,y)=N(x,y)-N(y,x)$. Table 1 shows a preference profile and its resulting preference matrix, $\mathbf{N}=(N(i,j))_{i,j}$, and margin matrix $\mathbf{M}=(\delta(i,j))_{i,j}=\mathbf{N}-\mathbf{N}^{T}$.
@@ 2.3 Social Choice Theory
The Condorcet winner defines a fairly intuitive concept: $A$ is the (strong) Condorcet winner if the number of votes where $A$ is ranked higher than $B$ is greater than vice versa for all other alternatives $B$. A weak Condorcet winner wins or ties in every head-to-head pairing. Formally, given a preference profile, []delimited-[]succeeds-or-equals[$\succeq$][ ⪰ ], a weak Condorcet winner is an alternative $a^{*}\in A$ such that $\forall a^{\prime}\in A,\delta(a^{*},a^{\prime})\geq 0$. If the inequality is strict for all pairs except $(a^{*},a^{*})$ then we call it a strong Condorcet winner. In the example shown in Table 1, alternative $C$ is the strong Condorcet winner. It dominates $A$ since three out of the five voters prefer $C$ to $A$. A similar situation holds when $C$ is compared to alternative $B$. While many have argued that this definition captures the essence of the correct collective choice (de Condorcet, 1785), in practice preference profiles may have no Condorcet winner. Condorcet-consistent voting schemes (e.g., Kemeny-Young introduced next) return a Condorcet winner when it exists, but differ on how they handle settings with no Condorcet winners.
@@
Table 1: Left: an example preference profile with five votes (Lanctot et al., 2023, Figure 1). The number on the left of the colon represents the number of votes of that type. Middle: the voter preference matrix, $\mathbf{N}$. Right: the voter margin matrix, $\mathbf{M}$.
Figure 1: An example calculation of the Kendall-tau distances of each vote in the preference profile from Table 1 to the optimal ranking $R=C\succ A\succ B$. The sum of the distances is 5. This sum for any other ranking $R^{\prime}\not=R$ is greater than 5.
2.3.1 Kemeny-Young Voting Method
The voting method was initially proposed by Kemeny (Kemeny, 1959). Later, its properties were characterized by Young & Levenglick  (Young and Levenglick, 1978). Let each ranking (total order) be represented as a permutation $\pi$ over $|A|$ alternatives. Define the Kemeny score of permutation $\pi$ as
$\textsc{KemenyScore}(\pi)=\sum_{(i,j),i<j}N(\pi[i],\pi[j]).$
@@ 2.3.1 Kemeny-Young Voting Method
The Kemeny rule returns a ranking that maximizes this: $\operatornamewithlimits{argmax}_{\pi\in\Pi(|A|)}\textsc{KemenyScore}(\pi)$. Kemeny-Young is Condorcet-consistent: if a Condorcet winner exists, it will be top-ranked by Kemeny-Young. It also satisfies the majority criterion, the Smith criterion (Smith, 1973), and monotonicity. The Kemeny rule always returns an optimal ranking, i.e., one whose sum of Kendall-tau distances to the votes is minimal, but its computational complexity is prohibitively expensive when $m$ is large.
2.4 Learning-to-Rank

### v3

@@ Abstract.
Driving progress of AI models and agents requires comparing their performance on standardized benchmarks; for general agents, individual performances must be aggregated across a potentially wide variety of different tasks. In this paper, we describe a novel ranking scheme inspired by social choice frameworks, called Soft Condorcet Optimization (SCO), to compute the optimal ranking of agents: the one that makes the fewest mistakes in predicting the agent comparisons in the evaluation data. This optimal ranking is the maximum likelihood estimate when evaluation data (which we view as votes) are interpreted as noisy samples from a ground truth ranking, a solution to Condorcet’s original voting system criteria. SCO ratings are maximal for Condorcet winners when they exist, which we show is not necessarily true for the classical rating system Elo. We propose three optimization algorithms to compute SCO ratings and evaluate their empirical performance. When serving as an approximation to the Kemeny-Young voting method, SCO rankings are on average 0 to 0.043 away from the optimal ranking in normalized Kendall-tau distance across 865 preference profiles from the PrefLib open ranking archive. In a simulated noisy tournament setting, SCO achieves accurate approximations to the ground truth ranking and the best among several baselines when 59% or more of the preference data is missing. Finally, SCO ranking provides the best approximation to the optimal ranking, measured on held-out test sets, in a problem containing 52,958 human players across 31,049 games of the classic seven-player game of Diplomacy.
Key words and phrases:
agent evaluation; social choice theory; ranking
1. Introduction
11footnotetext: Work performed while at Google DeepMind.
@@ 1. Introduction
Progress in the field of artificial intelligence has been driven by measuring the performance of agents on common benchmarks and challenge problems Samuel (1959); Campbell et al. (2002); Tesauro (1995); Syed ([n.d.]); Silver et al. (2016); Vinyals et al. (2019); (FAIR)† et al. (2022). In machine learning, common benchmarks like the UCI data set repository allowed direct comparisons of supervised learning algorithms Kelly et al. ([n.d.]). Competitions, such as ImageNet, led to breakthroughs in deep learning Krizhevsky et al. (2012).
@@
All of these examples require comparing agents (or models). Original success stories such as DeepBlue, TD-Gammon, and AlphaGo focused on a single domain. In the past ten years, agents have become increasingly more generally capable. AlphaZero extended application of AlphaGo to chess and Shogi Silver et al. (2018). The Arcade Learning Environment Bellemare et al. (2013), which steered much of the agent development in deep reinforcement learning, evaluated agents across 57 different Atari games. Recently, language models have been evaluated across suites of tasks such as in HELM Liang et al. (2022), BIG-bench bench authors (2023) AgentBench Liu et al. (2023), and via a public leaderboard such as Chatbot Arena driven by human voting Chiang et al. (2024). Answering simple questions for these generally capable agents, such as “Which is the best agent?” or “Is agent $X$ better than agent $Y$?” or “What is the relative ranking of agents $X$, $Y$, and $Z$?” become increasingly more difficult when aggregating over many different contexts: how agents are scored may vary wildly across tasks, data collected for evaluation may not be balanced evenly across tasks (or agents), and classical rating systems were simply not designed for this use case.
@@
To address these problems, recent ranking methods such as Vote’N’Rank Rofin et al. (2023) and Voting-as-Evaluation (VasE) Lanctot et al. (2023) use voting methods to aggregate results across tasks. Using computational social choice as a basis for ranking agents has several benefits: well-studied consistency properties of the voting methods are inherited, they do no require score normalization, and are less sensitive to score values and agent population than game-theoretic evaluation schemes Balduzzi et al. (2018). However, classic voting schemes, and related tournament solutions, typically assume that the data (e.g. agent comparisons) is complete. This assumption is not necessarily valid in the agent evaluation setting. While there has been research on identifying “necessary and possible winners” when there is incomplete voting or comparison data, the results are mixed Pini et al. (2011); Xia and Conitzer (2011); Aziz et al. (2015) and most of the findings are focused on identifying top-ranked agents, not ranking all agents as is our focus.
@@
In this paper, we introduce a new ranking scheme for general agents inspired by the interpretation of voting rules as maximum likelihood estimators Conitzer and Sandholm (2005). Starting with Condorcet’s original model of voting (Brandt et al., 2016, Chapter 8), Young showed that the maximum likelihood estimate (MLE) of the true ranking is the one that minimizes the sum of Kendall-tau distances to all the votes. Soft Condorcet Optimization (SCO) solves an optimization problem that assigns a numerical rating (score) $\theta_{a}$ to each agent (alternative) $a$. SCO then treats these ratings as a parameter vector, the votes as a data set, and defines a differentiable loss function as the objective, which can be optimized in several different ways. The final ranking of agents is obtained by sorting the ratings.
In summary, this paper makes the following contributions.
@@
1. SCO ranking scheme with the following properties: (1a) Three optimization methods to find ratings and corresponding rankings: gradient descent applied to a soft Kendall-tau distance (“sigmoid loss”), or a Fenchel-Young loss (perturbed optimization) Berthet et al. (2020); or solving a sigmoidal program with a branch-and-bound method Udell and Boyd (2014). (1b) Online form that can update ratings, and thus rankings, from individual outcomes as evaluation data arrives. (1c) Theorem 5 guarantees that the top-ranked agent by SCO ratings according to the sigmoid loss is the Condorcet winner when one exists.
2. Empirical evaluations that demonstrate the following: (2a) SCO ranking using sigmoid loss solves a failure mode of classical Elo rating system which may top-rank an agent that is not a Condorcet winner even when one exists. (2b) SCO can serve as an approximation to the Kemeny-Young voting method, indeed empirically finding low approximation error to the optimal ranking: on average 0 to 0.043 away in normalized Kendall-tau distance across 865 preference profiles from the PrefLib Mattei and Walsh (2013). (2c) In a noisy tournament setting with sparse data, SCO approximates the true ranking best when a large proportion (59% or more) of the data is missing. (2d) SCO ratings are closer to optimal rankings than Elo and voting-as-evaluation methods on held out test sets over 31,049 human Diplomacy games played by 52,958 players.
2. Background
@@
2.1. Evaluation of General Agents
We first paraphrase several key descriptions from Balduzzi et al. (2018); Lanctot et al. (2023). The problem of evaluating agents is that of ranking agents according to their skill. Skill can be determined in several ways. In the Agent-versus-Task (AvT) setting, agents compete individually in different tasks and compare outcomes (scores) to each other in each task (e.g., language models and the various metrics for assessing their abilities). In the Agent-versus-Agent (AvA) setting, agents directly compete against each other (e.g., online games such as chess or Diplomacy).
2.1.1. Classical Evaluation
@@
Elo has a number of positive qualities. First, it is a simple rule. Second, it can be used to estimate win rates between any two agents. Third, it can be easily employed online, i.e., to modify players’ ratings from the result of a single game. It is also a special case of logistic regression (for details, see Appendix D.2). Hence, Elo has been widely applied and is a common default choice for evaluation of agents. However, Elo has a number of well-known drawbacks Balduzzi et al. (2018); Lanctot et al. (2023); Bertrand et al. (2023). Of particular interest is the incompatibility of Elo with the concept of a Condorcet winner from social choice theory; examples are summarized in Section 4.1.
2.1.2. Voting as Evaluation of General Agents
Another way to evaluate agents is to use social choice theory, called Voting-as-Evaluation (VasE) recently proposed in Lanctot et al. (2023). In VasE, the alternatives correspond to general agents and votes to assessments of their performance. In the AvT setting, agents are ordered based on their performance on different tasks (such as each game in the Atari Learning Environment Bellemare et al. (2013) or on different benchmarks for large language models Liu et al. (2023)). In the AvA setting, agents compete directly in a multiagent environment, such as multiplayer games (like chess or poker), and each game outcome corresponds to a ranking over a subset of agents. Chatbot Arena Zheng et al. (2023), where language models compete head-to-head to provide the best answer to the same question, is another instance of the AvA setting. Casting agent evaluation as an application of computational social choice provides the benefit of robustness in the form of Condorcet consistency, clone/composition consistency, agenda consistency, and/or population consistency depending on the choice of voting rule used to evaluate agents. However, it is unclear how well these methods perform when the data is missing or unevenly distributed, which often occurs in the agent evaluation setting. SCO is motivated similarly to VasE and, as such, will be empirically assessed mainly for agent evaluation. To make the connection to ideas from the social choice literature, we will often use voting language when discussing SCO and its use in evaluation. Relevant terminology is introduced later in the paper.
2.2. Permutation and Ranking Distances
@@
2.3. Social Choice Theory
Figure 1. An example calculation of the Kendall-tau distances of each vote in the preference profile from Table 1 to the optimal ranking $R=C\succ A\succ B$. The sum of the distances is 5. This sum for any other ranking $R^{\prime}\not=R$ is greater than 5.
A voting scheme is defined as $\langle A,V,f\rangle$ where $A=\{a_{1},\ldots,a_{m}\}$ is the set of alternatives (agents), $V=\{v_{1},\ldots,v_{n}\}$ is the set of voters, and $f$ is the voting rule that determines how votes are aggregated. Voters have preferences over alternatives: $a_{1}\succ_{v_{i}}a_{2}$ indicates voter $v_{i}$ strictly prefers alternative $a_{1}$ over alternative $a_{2}$. In this paper, we assume strict preferences only, however the ideas can easily be extended to the case that allow weak preference (including ties/indifference between alternatives, e.g.$a_{1}\succeq s_{2}$). These preferences induce total orders over alternatives, which we denote by $\mathcal{L}$. A preference profile, $[\succ]\in\mathcal{L}^{n}$, is a vector specifying the preferences of each voter in $V$. It can be useful to summarize the preference profile in a voter preference matrix $\mathbf{N}$ or vote margin matrix $\mathbf{M}$. The preference count $N(x,y)$, $x,y\in A$ is the number of voters in []delimited-[]succeeds[$\succ$][ ≻ ] that strictly prefer $x$ to $y$. The vote margin is the difference in preference count: $\delta(x,y)=N(x,y)-N(y,x)$. Table 1 shows a preference profile and its resulting preference matrix, $\mathbf{N}=(N(i,j))_{i,j}$, and margin matrix $\mathbf{M}=(\delta(i,j))_{i,j}=\mathbf{N}-\mathbf{N}^{T}$.
@@ 2.3. Social Choice Theory
The Condorcet winner defines a fairly intuitive concept: $A$ is the (strong) Condorcet winner if the number of votes where $A$ is ranked higher than $B$ is greater than vice versa for all other alternatives $B$. A weak Condorcet winner wins or ties in every head-to-head pairing. Formally, given a preference profile, []delimited-[]succeeds[$\succ$][ ≻ ], a weak Condorcet winner is an alternative $a^{*}\in A$ such that $\forall a^{\prime}\in A,\delta(a^{*},a^{\prime})\geq 0$. If the inequality is strict for all pairs except $(a^{*},a^{*})$ then we call it a strong Condorcet winner. In the example shown in Table 1, alternative $C$ is the strong Condorcet winner. It dominates $A$ since three out of the five voters prefer $C$ to $A$. A similar situation holds when $C$ is compared to alternative $B$. While many have argued that this definition captures the essence of the correct collective choice (de Condorcet, 1785), in practice preference profiles may have no Condorcet winner. Condorcet-consistent voting schemes (e.g., Kemeny-Young introduced next) return a Condorcet winner when it exists, but differ on how they handle settings with no Condorcet winners.
@@
Table 1. Left: an example preference profile with five votes (Lanctot et al., 2023, Figure 1). The number on the left of the colon represents the number of votes of that type. Middle: the voter preference matrix, $\mathbf{N}$. Right: the voter margin matrix, $\mathbf{M}$.
2.3.1. Kemeny-Young Voting Method
@@
The voting method was initially proposed by Kemeny Kemeny (1959). Later, its properties were characterized by Young & Levenglick  Young and Levenglick (1978). Let each ranking (total order) be represented as a permutation $\pi$ over $|A|$ alternatives. Define the Kemeny score of permutation $\pi$ as $\textsc{KemenyScore}(\pi)=\sum_{(i,j),i<j}N(\pi[i],\pi[j])$. The Kemeny rule returns the ranking that maximizes this Kemeny score: $\operatornamewithlimits{argmax}_{\pi\in\Pi(|A|)}\textsc{KemenyScore}(\pi)$. Kemeny-Young is Condorcet-consistent: if a Condorcet winner exists, it will be top-ranked by Kemeny-Young. It also satisfies the majority criterion, the Smith criterion Smith (1973), and monotonicity. The Kemeny rule always returns an optimal ranking, i.e., one whose sum of Kendall-tau distances to the votes is minimal, but its computational complexity is prohibitively expensive when $m$ is large.
2.4. Learning-to-Rank